# Excited State Densities from Time-Dependent Density Functional Response Theory

### v1

@@ 3.1 One-Dimensional Helium: Local Excitations
Figure 1 shows the excited-state density differences of the first four local excitations of 1DHe. In this plot, the exact KS orbitals are used in Eq. 22 and $f_{\scriptscriptstyle\rm HXC}$ and $\tilde{g}_{\scriptscriptstyle\rm XC}$ are approximated by EXX: $f_{\scriptscriptstyle\rm HXC}(x,x^{\prime})=f_{\scriptscriptstyle\rm HX}(x,x^{\prime})=\frac{1}{2\sqrt{1+(x-x^{\prime})^{2}}}$, and $\tilde{g}_{\scriptscriptstyle\rm XC}=\tilde{g}_{\scriptscriptstyle\rm X}=0$. The plots show the fully converged SMA and the STL excited-state density differences, as well as the KS one; the inset shows the excited-state density itself, obtained from adding the exact ground-state density to the density-difference in each case.
@@
Figure 1: Excited-state density differences for 1D Helium: TDDFT within SMA, STL, compared with exact and KS, for the lowest 4 excitations (panels (a)–(d)). The insets show the exact excited-state density in each case. For the approximations, the exact KS orbitals are used in the formulas for the densities, and $f_{\scriptscriptstyle\rm XC}^{\rm EXX}$ is used for the kernel. Panel (e) shows the convergence for the lowest 3 excited densities with respect to the number of orbitals included in the sums over $p$.
@@
We see that SMA gives a noticeable correction to the KS excited-state densities, particularly for the lowest excitation, and that the SMA, STL, and KS densities all become more accurate for the higher excitations, as expected. In particular, for the lowest excitation, SMA captures the sharper peak and the shoulders around the location of the He nucleus, which are missed by the KS density in the lowest excitation, as seen in the inset. For this lowest excitation, the effect of the higher KS orbitals is important, since the STL density is almost indistinguishable from the KS density. For the higher excitations shown, STL brings a small correction to the KS density (in fact, almost none for the third excited state).
[…]
@@
In Figure 2, we compare the SMA densities predicted with the EXX and LDA approximations for the orbitals and the $f_{\scriptscriptstyle\rm HXC}$ kernel; we take again $\tilde{g}_{\scriptscriptstyle\rm XC}=\tilde{g}_{\scriptscriptstyle\rm X}=0$ throughout. We find that EXX performs notably better than LDA, particularly for the choice of the KS orbitals. Using $f_{\scriptscriptstyle\rm HXC}^{\rm LDA}$ with EXX orbitals tends to slightly under- and over-estimate the density difference near the position of He nucleus, but the effect of the kernel is less than the effect of the choice of the KS orbital approximation. Though SMA with LDA orbitals seems to estimate the density difference only slightly worse for the lowest excitation, for higher ones, it fails to capture the density difference accurately away from the nucleus, as might be expected since the too-rapid decay of the LDA KS potential causes significant errors in the KS orbitals away from the core.
@@ 3.2 Double-well potentials: Charge-Transfer Excitations
We begin with the double-well where the ground-state has both electrons in the soft-Coulomb potential on the left ($v_{\rm ext}^{\mathrm{gs-soft}}$), that had been studied in Ref. 33. The lowest excitation has a charge-transfer nature. In Fig. 3(a), we observe that, when exact KS orbitals are used to evaluate the sums in the expressions, the KS density is already a good approximation with only a small error to the density of the charge-transfer excitation, that STL gives practically no correction, while the SMA gives a larger correction, giving a density very close to the exact. The results in this figure used EXX for the kernel.
Figure 3: Excited-state density differences for the CT excitation of FERM13 double well model33. Panel (a) shows SMA, STL compared to the exact and KS excited-state density difference. Exact KS orbitals and $f_{\scriptscriptstyle\rm XC}^{\mathrm{EXX}}$ we used in approximations. The inset shows the excited state density of CT state. Panel (b) shows the performance of SMA with KS orbitals and $f_{\scriptscriptstyle\rm XC}$ approximated by LDA and EXX functionals. Panel (c) illustrates the issues of SMA and STL excited-state density differences when . Panel (d) shows the exact, EXX ground state densities, along with and LDA ground state which is wrongfully localized in the second well, and the corresponding KS potentials in the inset.
@@
Why does this problem emerge? LDA underestimates the gap between the occupied and lowest unoccupied KS levels: while the exact LUMO-HOMO gap is $\nu_{\mathrm{exact}}=0.112$ a.u., the LDA gap is far smaller, at $\nu_{\mathrm{LDA}}=0.005$ a.u. This means that for all excitations, the terms in Eq. (22) have much too small denominators, which in itself would lead to divergences unless cancelled out by a similar term in the numerator. For this lowest charge-transfer excitation, this cancellation does in fact occur, but the divergence remains because the gap appears also in the denominator of $\chi_{\scriptscriptstyle\rm S}$. When we approximate $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ by its first-order term in the Taylor expansion in $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$, the severely underestimated denominator leads to the divergence. Thus, the problem lies not with the SMA itself, but rather with our approximation of the inverse of $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ that appears in it.
@@
Instead, as discussed earlier, the STL approach resums the Taylor series (Eq. (23)) and so treats this inverse exactly, although within the STL approximation. In Fig. 3(c), we show that the STL density difference with LDA indeed does not display any divergence, while approximating the inverse function $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ by its first-order term in $f_{\scriptscriptstyle\rm HXC}$ in the same way as done in SMA does diverge. However, with LDA, the STL charge-transfer density-difference still gives substantial error compared to exact one, and this error stems from the erroneous LDA ground-state shown in Fig. 3(d). The LDA suffers from a delocalization error, and has an unphysical fractional charge in each well instead of asymptotically having a two-electron density fully on the left. The ground-state fractional charge error comes hand in hand with the vanishing LDA KS HOMO-LUMO gap, and this fractional-charge error is transferred to the KS charge-transfer excitation too. This is evident in Fig. 3(c) where integrating the KS curve on either side of zero falls short of 1. The wrong behavior of the KS orbitals leads to an erroneous STL density difference, even with the wrong sign.
@@
To emphasize the importance of an accurate KS ground-state, we perform the same calculations on the double-well model with the potential $v_{\rm ext}^{\mathrm{gs-loc}}$ from Ref. 36, where the two electrons in the ground-state are more tightly localized in the left well, so that LDA has a better hope to capture this qualitatively correctly. The second lowest excitation of this system is the charge transfer, the density of which is shown in Fig. 4.
@@
Figure 4: Excited-state density differences for the CT excitation in the double well model $v_{\rm ext}^{\mathrm{gs-loc}}$. Panel (a) shows SMA, STL compared to the exact and KS excited-state density difference, using the exact KS orbitals and $f_{\scriptscriptstyle\rm XC}^{\mathrm{EXX}}$. The inset shows the excited state density of CT state. Panel (b) shows the performance of SMA with KS orbitals and $f_{\scriptscriptstyle\rm XC}$ approximated by LDA and EXX functionals as indicated in the legend. Panel (c) illustrates the performance of STL compared to SMA and exact excited-stated density difference. Panel (d) shows the exact, EXX and LDA ground state densities, and corresponding KS potentials in the inset.
@@
The charge-transfer excited state density-differences obtained with exact KS orbitals and $f_{\scriptscriptstyle\rm HXC}^{\mathrm{EXX}}$ (Fig. 4(a)), as well as with EXX and LDA approximations (Fig. 4(b)), are all in good agreement with the exact. We highlight specifically the result given by LDA approximation, since unlike in the previous result, we not only do not encounter the divergences of the density difference, but it even agrees closely with the exact result. The KS excitation frequency is not qualitatively different from the exact, as it was in the previous case: the exact HOMO-LUMO gap is $\nu_{\mathrm{exact}}=2.235$ a.u., while LDA one is $\nu_{\mathrm{LDA}}=1.974$ a.u. We emphasize that getting reasonable KS quantities is of the most importance for the good performance of the excited state densities. Regarding the approximate evaluations of these densities, we observe that while both SMA and STL accurately capture the charge transfer in this case, STL performs slightly better. This can be explained by inspecting the behavior of LDA KS potential in the inset of Fig. 4(d): the lowest KS orbitals tend to be more accurate, leading to a good approximation of the ground state, but a worse approximation of excited-state density difference (second-excited state) with SMA.
@@ 3.3 Harmonic potentials: Double excitations
Fig. 5 shows the excited-state densities of the lowest three excitations, predicted from adiabatic SMA, SPA, frequency-dependent DSMA and DSPA, and compared with the exact. As expected, the approximations coincide closely with the exact for the lowest excitation, but differ for the second and third due to their double-excitation character. Both DSPA and DSMA do an excellent job, while the adiabatic SMA and SPA qualitatively fail for both states at $\gamma=0$. For the third state at $\gamma=1$ the adiabatic approximations also fail, but they provide a closer approximation for the 2nd state, with a small error at the origin, consistent with the fact that it is dominated by the single excitation. We note there is only one adiabatic state, so the predicted density for both the actual 2nd and 3rd states is the same with the adiabatic SMA/SPA.
Figure 5: Column 1 - 1st excitation of 1d Harmonic oscillator at degrees of mixing $\gamma=0$ and $\gamma=1$, Exact vs. SMA; Column 2 - 1st of the double excitations at $\gamma=0$ and $\gamma=1$, DSMS vs SMA vs Exact; Column 3 - 2nd of the double excitations at $\gamma=0$ and $\gamma=1$, DSMS vs SMA vs Exact.
@@
Thus, we have shown that the frequency-dependent kernels of Ref. 21, 23, which were shown earlier to produce good excitation energies for states of double-excitation character, and that of Ref. 23 also good oscillator strengths, both also yield excellent approximations to excited-state densities.

### v2

@@ 3.1 One-Dimensional Helium: Local Excitations
Figure 1 shows the excited-state density differences of the first four local excitations of 1DHe. In this plot, the exact KS orbitals are used in Eq. 22 and $f_{\scriptscriptstyle\rm HXC}$ and $\tilde{g}_{\scriptscriptstyle\rm XC}$ are approximated by EXX: $f_{\scriptscriptstyle\rm HXC}(x,x^{\prime})=f_{\scriptscriptstyle\rm HX}(x,x^{\prime})=\frac{1}{2\sqrt{1+(x-x^{\prime})^{2}}}$, and $\tilde{g}_{\scriptscriptstyle\rm XC}=\tilde{g}_{\scriptscriptstyle\rm X}=0$. The plots show the fully converged SMA and STL excited-state density differences, as well as the KS one; the inset shows the excited-state density itself, obtained from adding the exact ground-state density to the density-difference in each case.
@@
Figure 1: Excited-state density differences for 1D Helium: TDDFT within SMA, STL, compared with exact and KS, for the lowest 4 excitations (panels (a)–(d)). The insets show the total excited-state density in each case. For the approximations, the exact KS orbitals are used in the formulas for the densities, and $f_{\scriptscriptstyle\rm XC}^{\rm EXX}$ is used for the kernel. Panel (e) shows the convergence for the lowest 3 excited densities with respect to the number of orbitals included in the sums over pp.
@@
We see that SMA gives a noticeable correction to the KS excited-state densities, particularly for the lowest excitation, and that the SMA, STL, and KS densities all become more accurate for the higher excitations, as expected. In particular, for the lowest excitation, SMA captures the sharper peak and the shoulders around the location of the He nucleus, which are missed by the KS density in the lowest excitation, as seen in the inset. For this lowest excitation, the effect of the higher KS orbitals is important in the construction of $\chi_{\scriptscriptstyle\rm S}$ going into the formula, since the STL density is almost indistinguishable from the KS density. For the higher excitations shown, STL brings a small correction to the KS density (in fact, almost none for the third excited state).
Panel (e) shows the convergence of the SMA density-difference with the number $K$ of orbitals included in the sums over $p$ in Eq. 22, defined as $\sigma_{K}=\int dx\left|\Delta n_{I}^{K}(x)-\Delta n_{I}^{500}(x)\right|^{2}$, where $N$ is the number of grid points. We see that, for these states, even with just one orbital, this spatially-averaged error is relatively small, within 0.03 for the lowest excited state, and 0.01 for the 2nd-4th, and this error decreases rapidly with the number of orbitals, dropping to well within 0.00001 by 50 orbitals.
@@
In Figure 2, we compare the SMA densities predicted with the EXX and LDA approximations for the orbitals and the $f_{\scriptscriptstyle\rm HXC}$ kernel; in the cases when $f_{\scriptscriptstyle\rm HXC}^{LDA}$ is used, we also use the LDA approximation for $\tilde{g}_{\scriptscriptstyle\rm XC}=\tilde{g}_{\scriptscriptstyle\rm XC}^{LDA}$. We find that EXX performs notably better than LDA, particularly for the choice of the KS orbitals. Using $f_{\scriptscriptstyle\rm HXC}^{\rm LDA}$ with EXX orbitals tends to slightly under- and over-estimate the density difference near the position of He nucleus, but the effect of the kernel is less than the effect of the choice of the KS orbital approximation. Though SMA with LDA orbitals seems to estimate the density difference only slightly worse for the lowest excitation, for higher ones, it fails to capture the density difference accurately away from the nucleus, as might be expected since the too-rapid decay of the LDA KS potential causes significant errors in the KS orbitals away from the core.
@@ 3.2 Double-well potentials: Charge-Transfer Excitations
We begin with the double-well where the ground-state has both electrons in the soft-Coulomb potential on the left ($v_{\rm ext}^{\mathrm{gs-soft}}$), that had been studied in Ref. 43. The lowest excitation has a charge-transfer nature. In Fig. 3(a), we observe that, when exact KS orbitals are used to evaluate the sums in the expressions, the KS density is already a good approximation with only a small error to the density of the charge-transfer excitation, that STL gives practically no correction, while the SMA gives a larger correction, giving a density very close to the exact. The results in this figure used EXX for the kernel.
Figure 3: Excited-state density differences for the CT excitation in the double well model $v_{\rm ext}^{\mathrm{gs-soft}}$ 43. Panel (a) shows SMA, STL compared to the exact and KS excited-state density difference. Exact KS orbitals and $f_{\scriptscriptstyle\rm HX}$ were used in approximations. The inset shows the excited state density of CT state. Panel (b) shows the performance of SMA with KS orbitals and $f_{\scriptscriptstyle\rm HXC}$ approximated by LDA and EXX functionals. Panel (c) illustrates the issues of SMA and STL: “STL-approx” denotes the excited-state density difference when the first order expansion of $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ was used as an approximation, while STL with no such approximation shows no divergences. Panel (d) shows the exact, EXX ground state densities, along with and LDA ground state which has an incorrect delocalization to the right-hand well, and the corresponding KS potentials in the inset.
@@
Why does this problem emerge? LDA underestimates the gap between the occupied and lowest unoccupied KS levels: while the exact LUMO-HOMO gap is $\nu_{\mathrm{exact}}=0.112$ a.u., the LDA gap is far smaller, at $\nu_{\mathrm{LDA}}=0.005$ a.u. This means that for all excitations, the terms in Eq. 22 have much too small denominators, which in itself would lead to divergences unless cancelled out by a similar term in the numerator. For this lowest charge-transfer excitation, this cancellation does in fact occur, but the divergence remains because the gap appears also in the denominator of $\chi_{\scriptscriptstyle\rm S}$. As can be seen in Fig. 3(b), this yields a vastly overestimated density in the right well, with an integrated value to the right far larger than the value of 1 that is expected for the charge-transfer state. The density-difference is equally too negative in the left well, such that the total difference-density is zero (see Eq. 18 and discussion there), and the total density unphysically goes negative in the left region. When we approximate $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ by its first-order term in the Taylor expansion in $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$, the severely underestimated denominator leads to the divergence. Thus, the problem lies not with the SMA itself, but rather with our approximation of $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ that appears in it.
@@
Instead, as discussed earlier, the STL approach resums the Taylor series (Eq. 23) and so treats this inverse exactly, although within the STL approximation. In Fig. 3(c), we show that the STL density difference with LDA indeed does not display any divergence, while approximating the inverse function $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}$ by its first-order term in $f_{\scriptscriptstyle\rm HXC}$ in the same way as done in SMA does diverge. However, with LDA, the STL charge-transfer density-difference still gives substantial error compared to exact one, and this error stems from the erroneous LDA ground-state shown in Fig. 3(d). The LDA suffers from a delocalization error, and has an unphysical fractional charge in each well instead of asymptotically having a two-electron density fully on the left. The ground-state fractional charge error comes hand in hand with the vanishing LDA KS HOMO-LUMO gap, and this fractional-charge error is transferred to the KS charge-transfer excitation too. This is evident in Fig. 3(c) where integrating the KS curve on either side of zero falls short of 1. The wrong behavior of the KS orbitals leads to an erroneous STL density difference, even with the wrong sign.
@@
To emphasize the importance of an accurate KS ground-state, we perform the same calculations on the double-well model with the potential $v_{\rm ext}^{\mathrm{gs-loc}}$ from Ref. 46, where the two electrons in the ground-state are more tightly localized in the left well, so that LDA has a better hope to capture this qualitatively correctly. The second lowest excitation of this system is the charge transfer, the density of which is shown in Fig. 4.
@@
Figure 4: Excited-state density differences for the CT excitation in the double well model $v_{\rm ext}^{\mathrm{gs-loc}}$ 46. Panel (a) shows SMA, STL compared to the exact and KS excited-state density difference, using the exact KS orbitals and $f_{\scriptscriptstyle\rm HX}$. The inset shows the excited state density of CT state. Panel (b) shows the performance of SMA with KS orbitals and $f_{\scriptscriptstyle\rm HXC}$ approximated by LDA and EXX functionals as indicated in the legend. Panel (c) illustrates the performance of STL compared to SMA and exact excited-stated density difference. Panel (d) shows the exact, EXX and LDA ground state densities, and corresponding KS potentials in the inset.
@@
The charge-transfer excited state density-differences obtained with exact KS orbitals and $f_{\scriptscriptstyle\rm HX}$ (Fig. 4(a)), as well as with EXX and LDA approximations (Fig. 4(b)), are all in good agreement with the exact. We highlight specifically the result given by LDA approximation, since unlike in the previous result, we not only do not encounter the divergences of the density difference, but it even agrees closely with the exact result. The KS excitation frequency is not qualitatively different from the exact, as it was in the previous case: the exact HOMO-LUMO gap is $\nu_{\mathrm{exact}}=2.235$ a.u., while LDA one is $\nu_{\mathrm{LDA}}=1.974$ a.u. We emphasize that getting reasonable KS quantities is of the most importance for the good performance of the excited state densities. Regarding the approximate evaluations of these densities, we observe that while both SMA and STL accurately capture the charge transfer in this case, STL performs slightly better. This can be explained by inspecting the behavior of LDA KS potential in the inset of Fig. 4(d): the lowest KS orbitals tend to be more accurate, leading to a good approximation of the ground state, but a worse approximation of excited-state density difference (second-excited state) with SMA.
@@ 3.3 Harmonic potentials: Double excitations
Fig. 5 shows the excited-state densities of the lowest three excitations, predicted from adiabatic SMA, SPA, frequency-dependent DSMA and DSPA, and compared with the exact. The results in Fig. 5 utilize the exact KS orbitals and exact KS orbital energies in the construction of Eqs. (22), (31), (38), together with the EXX kernel for the adiabatic part. As expected, the approximations coincide closely with the exact for the lowest excitation, but differ for the second and third due to their double-excitation character. Both DSPA and DSMA do an excellent job, while the adiabatic SMA and SPA qualitatively fail for both states at $\gamma=0$. We note the adiabatic approximation predicts only one state, so the predicted density for both the actual 2nd and 3rd states is the same with the adiabatic SMA/SPA. For the third state at $\gamma=1$ the adiabatic approximations also fail, but they provide a closer approximation for the second state, with a small error at the origin, consistent with the fact that it is dominated by the single excitation. We observe that for DSPA and DSMA, the second excitation performs worse compared to ASPA/ASMA, and we will address that shortly when discussing the performance of LDA kernel utilized with the exact KS orbitals in the same setting.
Figure 5: Excited state densities of the Harmγ model with the set of exact orbitals and EXX kernel: Panel (a) First excitation with $\gamma=0$, showing close agreement of exact, adiabatic SMA and SPA densities; Panels (b) and (c), Second and third excitations showing close agreement of DSMA and DSPA with exact, and larger errors of adiabatic SMA and SPA, due to the missing double-excitation character of the latter (see text); Panel (d) Lowest excitation with $\gamma=1$, again showing close agreement of exact, adiabatic SMA and SPA densities; Panels (e) and (f), Second and third excitations with $\gamma=1$ where ASMA outperforms DSMA and DSPA for the second excitation, however, DSMA and DSPA still are in a close agreement with the exact result.
In Fig. 6, we apply the LDA kernel for $f_{\scriptscriptstyle\rm HXC}^{A}$ of Eqs. (22), (31), (38), and immediately observe an increased difference between ASPA and ASMA densities, as well as between DPSA and DSMA ones. Upon inspecting the different terms, we find that most of the difference between the ASPA and ASMA appears to be due to the $g_{\scriptscriptstyle\rm XC}^{LDA}$ term and this is due to a factor of two difference in the $\chi_{\scriptscriptstyle\rm S}$: in the SPA Eq. (39), for consistency, we use the response function in the Tamm-Dancoff approximation, which in the static limit is different from the response function Eq. (40) used in Eq. (22) by the factor 1/2, that is,
$\chi_{\scriptscriptstyle\rm S}^{\mathrm{TDA}}(r,r{{}^{\prime}},\omega=0)=-2\sum_{a}^{\mathrm{unocc}}\frac{\phi_{0}(x)\phi_{a}(x)\phi_{a}(x{{}^{\prime}})\phi_{0}(x{{}^{\prime}})}{\epsilon_{a}-\epsilon_{0}}.$ (41)
Note that in the previous figure, where $f_{\scriptscriptstyle\rm HX}$ was utilized, $g_{\scriptscriptstyle\rm XC}=0$, so this difference did not arise. Differences in ASMA and ASPA densities contribute to DSMA and DSPA density-differences.
Interestingly, unlike with the EXX kernel, DSMA with the LDA kernel performs a little worse than DSPA for the second excitation at the limit $\gamma=0$ of Harmγ model, overestimating the density at the origin and underestimating it in the vicinity, while in the limit of $\gamma=1$, it performs best compared to all other approximations. In the latter case, ASPA density outperforms the ASMA one, however, this does not translate to an advantage of DSPA over DSMA. For the third excitation, even though DSPA exhibits an error at the origin, both dressed approximations perform well compared to the exact density.
In Figure 7, we show the DSMA doubly-excited state density diffrences obtained with the sets of EXX and LDA orbitals used in Eq. 31 together with EXX and LDA for $f_{\scriptscriptstyle\rm HXC}^{A}$. Since both provide accurate approximations for the most relevant KS orbitals, most of the differences in the resulting densities are mostly due to the $f_{\scriptscriptstyle\rm HXC}^{A}$ kernel. Therefore, the results obtained do not differ much compared to those shown in Fig. 5 for the densities calculated with EXX kernel, and to Fig. 6 for those calculated with LDA kernel, both on top of the exact KS orbitals and energies.
Figure 6: Excited state densities of the Harmγ model with the set of exact orbitals and LDA kernel: Panel (a) First excitation with $\gamma=0$, where ASPA outperforms ASMA but still exhibiting error at the origin; Panels (b) and (c), Second and third excitations showing closer agreement of DSPA with exact over DSMA in the second exciatiotion, and vice-versa in the third excitation, with qualitatively larger errors of adiabatic SMA and SPA, due to their missing double-excitation character; Panel (d) Lowest excitation with $\gamma=1$, again showing closer agreement of ASPA density with exact over ASMA; Panels (e) and (f), Second and third excitations with $\gamma=1$ where DSMA outperforms both adiabatic approximations and DSPA in the second excitation, while in the third, both DSPA and DSMA are in a close agreement with the exact result.
Table 1 shows the energies and approximate single excitation character $G_{I}^{2}$. The (slightly) larger error in the DSMA predictions of the second excited-state density when $f_{\scriptscriptstyle\rm HXC}^{LDA}$ is used for the ingredients in the DSMA in Fig. 6(b) and Fig. 7(a) is consistent with its larger errors in the excitation energy for the lower of the two states of double-excitation character in the $\gamma=0$ case. Yet, the accuracy of the excitation frequency under $f_{\scriptscriptstyle\rm HX}$ does not guarantee a better prediction of the corresponding density for $\gamma=1$, as seen in Fig. 5(e) and Fig. 7(c).
$\gamma$ Approximation $\omega^{\text{adia}}$ $\omega_{2}^{{DSPA}}$ $\omega_{2}^{{DSMA}}$ $\omega_{3}^{{DSPA}}$ $\omega_{3}^{{DSMA}}$ $G_{2}^{2}$ $G_{3}^{2}$
$\gamma=0$ Exact 1.73 1.73 2.00 2.00 0.56 0.44
$\{\phi^{\text{exact}}\}+f_{\scriptscriptstyle\rm HX}$ 1.86 1.72 1.72 2.01 2.01 0.52 0.48
$\{\phi^{\text{exact}}\}+f_{\scriptscriptstyle\rm HXC}^{{LDA}}$ 1.83 1.70 1.70 1.99 1.99 0.56 0.44
$\{\phi^{EXX}\}+f_{\scriptscriptstyle\rm HX}$ 1.87 1.72 1.72 2.01 2.01 0.50 0.50
$\{\phi^{{LDA}}\}+f_{\scriptscriptstyle\rm HXC}^{{LDA}}$ 1.83 1.70 1.70 1.99 1.99 0.57 0.43
$\{\phi^{EXX}\}+f_{\scriptscriptstyle\rm HXC}^{{LDA}}$ 1.84 1.71 1.71 2.00 2.00 0.54 0.46
$\{\phi^{{LDA}}\}+f_{\scriptscriptstyle\rm HX}$ 1.85 1.71 1.72 2.01 2.01 0.52 0.48
$\gamma=1$ Exact 2.60 2.60 2.98 2.98 0.88 0.12
$\{\phi^{\text{exact}}\}+f_{\scriptscriptstyle\rm HX}$ 2.66 2.61 2.61 2.99 2.99 0.85 0.15
$\{\phi^{\text{exact}}\}+f_{\scriptscriptstyle\rm HXC}^{{LDA}}$ 2.63 2.57 2.57 2.98 2.98 0.88 0.12
$\{\phi^{EXX}\}+f_{\scriptscriptstyle\rm HX}$ 2.67 2.62 2.61 2.99 2.99 0.85 0.15
$\{\phi^{{LDA}}\}+f_{\scriptscriptstyle\rm HXC}^{{LDA}}$ 2.63 2.58 2.58 2.98 2.98 0.87 0.13
$\{\phi^{EXX}\}+f_{\scriptscriptstyle\rm HXC}^{{LDA}}$ 2.63 2.58 2.58 2.98 2.98 0.87 0.13
$\{\phi^{{LDA}}\}+f_{\scriptscriptstyle\rm HX}$ 2.66 2.61 2.61 2.99 2.99 0.85 0.15
Table 1: Comparison of adiabatic ASMA, DSMA and DSPA frequencies, and DSMA single-excitation character $G_{I}^{2}$ using different approximations for the orbitals and kernels in Eq. 31.
Figure 7: Excited-state density differences for the second and the third excitations of Harmγ model: comparison of approximations for KS orbitals and $f_{\scriptscriptstyle\rm HXC}$ within DSMA. Panel (a) shows the better agreement of the exact and DSMA densities with EXX kernel over the DSMA with LDA kernel in the second excitation for $\gamma=0$, while Panel (c), on the contrary, shows better performance of LDA kernel in the second excitation for $\gamma=1$; Panel (b) and Panel (d) show close agreement of DSMA with both $f_{\scriptscriptstyle\rm HXC}$ in both limits of $\gamma$.
@@
Overall, we have shown that the frequency-dependent kernels of Ref. 31, 33, which were shown earlier to produce good excitation energies for states of double-excitation character, and that of Ref. 33 also good oscillator strengths, both also yield excellent approximations to excited-state densities. We note that the behavior of the densities under EXX and LDA kernels used for adiabatic part in DSMA is not always consistent with their relative performance for the energies of these states given by the same approach. To understand the underlying reasons of such behavior, further exploration is required. Still, the differences are very small compared to the improvement from the adiabatic description which entirely misses one state.